A hollow sphere rolls without sliding on an incline with it’s lowermost part touching the incline moving with velocity V,if we coat it with wax it slide down with velocity (5V)/4. find radius of gyration?
 

Grade:12th pass

1 Answers

Kenny Toijam
7 years ago
Radius of Gyration will be 3/4 R. For pure rolling we can easily derive that mgH=1/2mv2{1+(K/R)^2}.....................(1)For sliding after waxing we can applymgH=1/2 m(5v/4)^2...............................(2)Equating equation 1 and 2 we can conclude that K=(3/4)RHope it helps! -Kenny Toijam